OpenStudy (anonymous):

Solve the equation. 3(4x - 3) - 19 = 8x - 4

OpenStudy (anonymous):

multiply through and subtract like terms

OpenStudy (anonymous):

3(4x - 3) - 19 = 8x - 4 distribute 12x - 9 - 19 = 8x - 4 combine like terms 12x - 28 = 8x - 4 add four to both sides 12x - 24 = 8x subtract 12x from both sides -24 = -4x divide by -4 6 = x

hero (hero):

3(4x - 3) - 19 = 8x - 4 12x-9 -19 = 8x-4 12x - 28 = 8x-4 12x-8x = -4 + 28 4x = 24 x = 24/4 x = 6